How do you answer the following questions?
Given f(x)= x – 3 and g(x)=square root of (x-2)find the domain of the quotient function f/g.

Answer by stanbon(75887)

How do you answer the following questions?
Given f(x)= x – 3 and g(x)=square root of (x-2)find the domain of the quotient function f/g.
---------------
f/g = (x-3)/(sqrt(x-2))
Domain:
The "x-2" must not be negative.
x-2 >= 0
x >=2
---------
Domain: All Real Numbers x, where x>=2.
========================
Cheers,
Stan H.